OpenStudy (anonymous):

Solve for x: logbase4((x/2) + 4) = 3 Can someone explain this to me?

OpenStudy (helder_edwin):

\[ \large \log_4(x/2+4)=3 \]

OpenStudy (anonymous):

right

OpenStudy (helder_edwin):

use this \[ \large \log_ab=y\Leftrightarrow a^y=x \]

OpenStudy (anonymous):

4^3 = (x/2 + 4)?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why do you do that?

OpenStudy (helder_edwin):

now add -4 to both sides

OpenStudy (helder_edwin):

do what?

OpenStudy (anonymous):

what is "log", I can't remember

OpenStudy (helder_edwin):

log = logarithm

OpenStudy (anonymous):

x = 120

OpenStudy (anonymous):

what is a logarithm

OpenStudy (helder_edwin):

yes. x=120

OpenStudy (anonymous):

I remember what a logarithm is now. It's weird, I've forgotten a lot of stuff since last year. That inverse question you answered for me a minute ago would've been a breeze before. I don't know why I forget

OpenStudy (helder_edwin):

when u don't practice, u forget. the same happens to me with several var. calc.